the kinetic energy depends on speed square error in speed is 30 percent find error percentage in kinetic energy

Grade:11

1 Answers

Arun
4 years ago

k.e= m*(v^2)/2.
ln(k.e)=ln(m)+ 2ln(v)- ln(2).
applying differentiation on both sides and multiplying by 100 on both sides, gives the percentage error in k.e= 2*(percentage error in velocity)

hence answer is 60%